OpenStudy (anonymous):

Determine the number of solutions for each linear system in two-space. If possible, solve each system. a) 2x -5y = 6 {x = -9 + 7t {y = -4 + 3t

OpenStudy (anonymous):

2*(-9+7t) -5*(-4+3t)=6 -18+14t+20-15t=6 t=-4, x=-37, y=-16

OpenStudy (anonymous):

or is this trickier? :-)